# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v1

@@ 4.5 Discussion
Our numerical experiments provide insights into different market clearing models, pricing rules, and their economic implications on day-ahead markets. Note that the final dispatch is affected by additional factors such as intraday trading or long-term capacity allocations. Nodal pricing tends to result in slightly higher average prices compared to zonal pricing, irrespective of the chosen zonal configuration. However, Transmission constraints can be violated and necessitate costly redispatch measures. As a result, a zonal pricing will always pick an inefficient dispatch, and even if redispatch is conducted at minimal costs, the total system costs will be substantially higher than under nodal pricing. Given the steep increase of redispatch costs in recent years, this effect is likely to magnify in the future.
Interestingly, the choice of zone configuration did not substantially impact prices. All zonal configurations suggested in the Bidding Zone Review simplify the network to an extent that opting for two, three, or four zones yields similar results regarding the average prices and the price standard deviation. In contrast to nodal pricing, zonal pricing leads to substantial redispatch and higher total costs. While nodal prices may vary between nodes, these differences incentivize short-term demand response on nodes where this is most helpful for overall system stability. If such price differences persist over more extended time periods, they also set investment incentives for generators and grid expansion.
Given the absence of Walrasian equilibria, any pricing rule – regardless of zonal or nodal pricing – must balance economic properties such as welfare gains and the participants’ lost opportunity costs. The currently used Euphemia algorithm leads to welfare losses and is computationally expensive but avoids paradoxically accepted bids. Our experiments suggest that non-uniform pricing rules achieve low MWPs for paradoxically accepted bids while maintaining the optimal allocation and scaling in polynomial time. Our findings thus provide support for non-uniform pricing as they are currently being discussed in the context of the Capacity Allocation and Congestion Management (CACM) Regulations (All NEMO Committee, 2023b, ).
CH pricing minimizes GLOCs, but our results confirm previous findings that congestion signals are distorted (evidenced by high LLOCs) and high MWPs are necessary. IP pricing, the prevailing rule in U.S. markets, ensures effective congestion signals but requires very high MWPs. The Join pricing rule strikes a balance between LLOCs and MWPs, ensuring low side-payments and good congestion signals. GLOCs are less of a concern, because already today, in U.S. markets with IP pricing, such incentives to deviate from the computed dispatch are mitigated by penalties for deviations.
@@ 5 Conclusion
The discussion surrounding zonal and nodal pricing has a long history. The European electricity market has employed zonal pricing for many years, often with large, nationwide price zones. However, the ongoing energy transition requires a reevaluation of price zones. Based on recent data released by ENTSO-E in the context of the EU Bidding Zone Review process, we compare various allocation and pricing rules for the German power market. Our findings indicate that, in terms of system costs and price signals, a nodal allocation rule with non-uniform pricing leads to the lowest total costs. Costly redispatch is mitigated, and the average prices increase only slightly. Additionally, pricing rules such as Join pricing require very low side-payments, while efficiently signaling transmission bottlenecks. In contrast, the proposed zonal configurations only differ marginally regarding system costs and the differences in average prices across zones are low. Interestingly, the effect of zonal splits on price standard deviations is low, and we do not find evidence for significant reductions in redispatch costs resulting from such splits, all other factors held constant.
@@ Acknowledgements
We would like to thank Karsten Neuhoff and Maximilian Rinck for many helfpul comments. This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. BI 1057/10.

### v4

@@ 4.5 Discussion
Our numerical experiments with the BZR data set provide insights into different market clearing models, pricing rules, and their short-term economic implications on day-ahead markets. Nodal pricing tends to result in slightly higher average prices compared to zonal pricing, irrespective of the chosen zonal configuration. However, transmission constraints can be violated and necessitate costly redispatch measures. As a result, a zonal pricing selects an inefficient dispatch, and even if redispatch is conducted at minimal costs, the total system costs will be significantly higher than under nodal pricing. Given the steep increase of redispatch costs in recent years, this effect is likely to magnify in the future.
Interestingly, the choice of zonal configuration did not substantially impact prices. All zonal configurations suggested by ACER in the BZR simplify the network to an extent that opting for two, three, or four zones yields similar results regarding the average prices and the price standard deviation. Analyses of individual days show that congestion arises, but not necessarily at the cross-zonal links but within zones. While nodal prices may vary between nodes, these differences incentivize short-term demand response on nodes where this is most helpful for overall system stability. If such price differences persist over more extended time periods, they also set investment incentives for generators and grid expansion. Note that our analysis does not take such incentives into account and assumes static demand. Taking incentives for demand reduction and targeted investments into account would lead to more similar prices rather in the long run.
Given the absence of Walrasian equilibria, any pricing rule – regardless of zonal or nodal pricing – must balance economic properties such as welfare gains and the participants’ lost opportunity costs. The currently used Euphemia algorithm leads to welfare losses and is computationally expensive but avoids paradoxically accepted bids. Our experiments suggest that non-uniform pricing rules achieve low make-whole payments for paradoxically accepted bids while maintaining the optimal allocation and scaling in polynomial time. The findings provide support for non-uniform pricing as it is currently being discussed in the context of the Capacity Allocation and Congestion Management (CACM) Regulations (All NEMO Committee,, 2023). CH pricing minimizes GLOCs, but our results confirm previous findings that high MWPs are necessary. IP pricing, the prevailing rule in U.S. markets, ensures effective congestion signals but requires very high MWPs. The Join pricing rule strikes a balance between LLOCs and MWPs, ensuring low side-payments and good congestion signals.
@@ 5 Conclusion
The discussion surrounding zonal and nodal pricing has a long history. The European electricity market has employed zonal pricing for many years, often with large, nationwide price zones. However, the ongoing energy transition requires a reevaluation of price zones. Based on recent data released by ENTSO-E in the context of the EU Bidding Zone Review process, we compare various zonal and nodal pricing rules for the German power market. Our findings indicate that, in terms of system costs and price signals, a nodal allocation rule with non-uniform pricing leads to the lowest total costs. Costly redispatch is mitigated, and the average prices increase only slightly. Additionally, pricing rules such as Join pricing require very low side-payments, while efficiently signaling transmission bottlenecks. In contrast, the proposed zonal configurations only differ marginally regarding system costs and the differences in average prices across zones are low. Also, the effect of zonal splits on price standard deviations is low, and we do not find evidence for significant reductions in redispatch costs resulting from such splits. The study does not say that there cannot be congestion on cross-zonal links leading to significant price differences between zones at certain times. However, in the BZR dataset that was intended to inform ACER based on realistic demand and supply scenarios (the 2009 climate year), we do not find significant price differences. Fears of substantial price differences after a zonal split might be overrated based on this comprehensive and up-to-date dataset. Importantly, we could provide a fair comparison with nodal pricing rules and show that the welfare gains are around 9-10%, depending on the redispatch model.
@@ Acknowledgements
We would like to thank Karsten Neuhoff, Maximilian Rinck, and Anthony Papavasiliou for helfpul comments on earlier versions of the manuscript. Mistakes are of course all ours. This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. BI 1057/10.
Appendix A Matching of Generators to Grid Locations
The BZR grid data specify the nominal capacity of thermal generating units at each node, but information on the specific plant type and its operating restrictions are not provided. A different data set provides the aggregate capacity of each specific plant type (e.g., Hard Coal old, Hard Coal new, Gas CCGT, etc.) as well as their operating characteristics (e.g., minimum uptime constraints). Unfortunately, no straightforward mapping exists between the two datasets.
As discussed in Section 3, we first identify each thermal unit’s broad plant type (e.g., hard coal, gas, etc.) by matching IDs with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur. Next, we need to distribute these plants to the sub-categories provided in the seller data. For example, all hard coal units must be distributed between Hard Coal old1, Hard Coal old2, and Hard Coal new, with each sub-category having different operational characteristics.
To perform this mapping, we solve an integer program (Schmitt,, 2023). Let $K$ be the set of broad plant types (e.g., hard coal, gas) and $A_{k}$ be the set of sub-categories (e.g., Hard Coal old1, Hard Coal old2, Hard Coal new) for each type $k$. The total capacity and number of units of each specific plant type $a\in A_{k}$ is given as $P_{a}$ and $n_{a}$ from the seller data. From the grid data, we obtain a set of generators $I_{k}$ for each broad plant type $K$, each having a nominal capacity of $p_{i},i\in I_{k}$. We denote as $x_{ia}\in\{0,1\}$ as the desired mapping of each generator to a specific plant type.
With this notation, the mapping problem looks as follows:
$\min_{x}\quad$ $\sum_{k\in K,a\in A_{k}}-n_{a}(P_{a}-\sum_{i\in I_{k}}p_{i}x_{ia})+\sum_{k\in K,a\in A_{k}}-P_{a}(n_{a}-\sum_{i\in I_{k}}x_{ia})$ (9)
subject to $\sum_{a\in A_{k}}x_{ia}=1\quad\forall i\in I_{k},k\in K$
$0\leq P_{a}-\sum_{i\in I_{k}}p_{i}x_{ia}\geq 600\quad\forall a\in A_{k},k\in K$
$0\leq N_{a}-\sum_{i\in I_{k}}x_{ia}\geq 2\quad\forall a\in A_{k},k\in K$
The objective function minimizes the aggregate differences in both capacities and number of units per plant type, with equal weight. The first constraint ensures that each unit of broad type $k$ (e.g., hard coal) is assigned to exactly one specific plant type $a\in A_{k}$ (e.g., Hard Coal old1). The last two constraints bound the maximum allowed deviations in capacities and the number of units for each specific plant type $a$. Note that a broad type $k$ could not be identified for all units, and such unidentified units could be matched freely to any specific type $a$. After solving the optimization problem, we obtain a specific type $a$ for each unit in the grid model and can use its respective operating characteristics as input to the clearing problem in (1).
The result of this mapping is presented in Table 13. Especially hard coal, lignite, and oil plants could be mapped almost perfectly. Concerning gas units, the grid and seller datasets exhibited more significant differences, which is reflected in the mapping.
Seller Data Grid Data
Fuel / PlantType Capacity [MW] Units Capacity [MW] Units
Hard Coal / old1 1,905.1 8 1,905 8
Hard Coal / old2 5,172.4 27 5,165.4 27
Hard Coal / new 5,178 8 5,178 7
Lignite / old1 4,989.3 13 4,989 13
Lignite / old2 3,111.1 11 3,111 11
Lignite / new 6,437.8 11 6,437 9
Gas Conventional / old1 776.01 26 775.97 26
Gas Conventional / old2 1215.4 23 1,215.33 23
Gas CCGT / old1 3,601.049 78 3,001.054 78
Gas CCGT / old2 5,356.22 40 5,055.723 38
Gas CCGT / new 7,389.1 45 6,789.321 43
Gas OCGT / old 760.1 15 760.025 15
Gas OCGT / new 2,613.168 66 2,013.224 66
Gas CCGT / present1 64 2 64 2
Gas CCGT / present2 282 2 282 2
Light oil 998.33 21 997.43 21
Table 13: Generator Mapping Results
Appendix B Example of Non-Decreasing Redispatch Costs
The following hypothetical example shows that a configuration with $k$ zones does not necessarily lead to lower redispatch costs compared to a configuration with less than $k$ zones.
Example 1.
Consider an electricity market with sellers $S=\{s_{1},s_{2},s_{3},s_{4}\}$, a buyer $b_{1}$ and an underlying transmission network $G=(V,E)$ with nodes $V=\{v_{1},v_{2},v_{3},v_{4},v_{5},v_{6}\}$ and lines $E=\{\{v_{3},v_{5}\},\{v_{4},v_{5}\}\}$. The line capacities are 50 MW for $\{v_{3},v_{5}\}$ and 100 MW for $\{v_{4},v_{5}\}$, and the two lines have the same susceptance value. The buyer $b_{1}$ is located in node $v_{5}$ and has an inelastic demand of 100 MWh. Table 14 shows, for each seller, the node in which they are located and their generation cost. Assume that each seller can generate 200 MWh. For this electricity market, consider the two zonal configurations described in Table 15.
Seller Node Cost (EUR/MWh)
$s_{1}$ $v_{1}$ 1
$s_{2}$ $v_{2}$ 3
$s_{3}$ $v_{3}$ 2
$s_{4}$ $v_{4}$ 40
Table 14: Example - Characteristics of the sellers
Configuration Zone 1 Zone 2 Zone 3 Generation Redispatch
2 Zones $\{v_{1},v_{2},v_{3},v_{4}\}$$\{v_{5},v_{6}\}$ 100 2,200
3 Zones $\{v_{2},v_{4},v_{5}\}$$\{v_{1},v_{3}\}$$\{v_{6}\}$ 200 2,300
Table 15: Example - Zonal configurations. Generation and redispatch costs are expressed in EUR.
In the 2-zone configuration, the cross-zonal capacity is equal to 150 MW. In zonal market clearing, seller $s_{1}$ generates 100 MWh to fulfill the demand of $b_{1}$. This solution is not physically feasible. The units generated by $s_{1}$ cannot be delivered to $b_{1}$ since $v_{1}$ and $v_{5}$ are not connected. A feasible solution is obtained using the minimum-cost redispatch method described in Section 2.2.2. In this solution, $s_{3}$ and $s_{4}$ each produce 50 MWh, resulting in redispatch costs equal to $1\cdot 100+2\cdot 50+40\cdot 50=2,200$ EUR.
In the 3-zone configuration, the capacity between zones 1 and 2 is 50 MW. In zonal market clearing, sellers $s_{1}$ and $s_{2}$ generate 50 MWh each. This solution is not physically feasible, since $v_{1}$ and $v_{2}$ are not connected to $v_{5}$, hence the units generated by $s_{1}$ and $s_{2}$ cannot be delivered to $b_{1}$. The redispatch method computes a feasible solution in which $s_{3}$ and $s_{4}$ produce 50 MWh each. The redispatch costs are $1\cdot 50+2\cdot 50+3\cdot 50+40\cdot 50=2,300$ EUR.